9. Assume that you wish to purchase a 20-year bond that has a maturity value of $1,000 and

makes semiannual interest payments of $40. If you require a 10 percent effective annual return

on your investment, what is the maximum price you should be willing to pay for the bond?

(pick the closest number) answer 846 how to get that
Answer #1

First, convert the effective annual rate to semi-annual APR as follows:

=((1+0.1)^(1/2))-1

=4.88%

Number of payments = 20*2 =40

In excel the formula is:

=PV(rate,nper,pmt,fv)

=PV(4.88%,40,40,1000)

=846

Also, in financial calculator, solve for PV, by entering I/Y 4.88%, N=40,PMT=40,FV=1000

Don't forget to upvote, if it was helpful, thank you!